Maritza is 3 years less than 4 times as old as her daughter. If the sum of their age is 37, how old is the mother?
Answer by ptaylor(2198)

Let x=daughter's age
Then 4x-3=Maritza's age
And we are told that x+4x-3=37
add 3 to each side
x+4x-3+3=37+3 Collect like terms
5x=40
x=8 daughter's age
4x-3=32-3=29 Maritza's age
CK
29+8=37
37=37
Hope this helps--ptaylor